# “Whose Side Are You On?” Estimating Ideology of Political and News Content Using Large Language Models and Few-shot Demonstration Selection

### v2

@@ Appendix E Sources can be misleading
Table 9 reports the accuracy of the LLMs on these specific test cases for liberal and conservative sources respectively. We see that the LLMs suffer from incorrect classifications when only the source is considered as the ground truth contradicts the ideology of the source. Here, providing the title increases the accuracy significantly especially in the conservative sources where the accuracy on just the source is for the YouTube and Ad Fontes datasets. In fact, providing the title for the YouTube test set in this case still does not improve the accuracy seeing as the model accuracy is still . This suggests that the LLM overtly relies on the source in this case completely disregarding title wherein accuracy on just the title was close to . Similar trends can be seen for the Ad Fontes news test set where the accuracy of source alone was and title was , and both was . We observe that the AllSides news test set still has better performance when the source is provided at a . On closer inspection, we observe that in the test set sampled, there is a high correlation between the source and article ideology which likely leads to such a high performance when source is provided.
Appendix F Statistical significances
We use the McNemar test to determine the statistical differences between the accuracies of two specific classifiers. Tables 10, 11, and 12 show the significance values for the models trained on the YouTube Slant, Ad Fontes news, and AllSides news datasets respectively.
| GPT | Llama | Mistral | ||||||||||||||||||||
| T | TS | TB | TSB | S | B | SB | T | TS | TB | TSB | S | B | SB | T | TS | TB | TSB | S | B | SB | ||
| GPT | T | — | 0.13 | 0.06 | 33.06** | 0.05 | 23.58** | 8.77** | 34.18** | 0.21 | 11.05** | 10.75** | 2.56 | 39.32** | 62.01** | 4.50* | 4.21* | 2.34 | 0.39 | 1.46 | 46.89** | 15.42** |
| TS | 0.13 | — | 0.52 | 27.95** | 0.52 | 30.99** | 13.77** | 35.79** | 0.69 | 12.10** | 12.30** | 4.69* | 41.32** | 69.72** | 4.72* | 2.79 | 3.31 | 0.04 | 2.77 | 53.35** | 20.66** | |
| TB | 0.06 | 0.52 | — | 42.29** | 0.00 | 28.96** | 10.27** | 30.30** | 0.06 | 10.25** | 10.16** | 2.13 | 37.80** | 65.91** | 2.96 | 5.33* | 1.70 | 0.88 | 1.07 | 54.30** | 18.06** | |
| TSB | 33.06** | 27.95** | 42.29** | — | 25.95** | 89.19** | 67.10** | 89.28** | 22.45** | 59.74** | 58.72** | 33.34** | 96.75** | 132.08** | 41.44** | 7.87** | 45.75** | 22.90** | 29.36** | 119.71** | 72.90** | |
| S | 0.05 | 0.52 | 0.00 | 25.95** | — | 18.40** | 8.13** | 28.34** | 0.03 | 7.84** | 8.07** | 3.54 | 34.04** | 65.24** | 2.02 | 5.47* | 0.98 | 0.76 | 1.90 | 38.75** | 12.04** | |
| B | 23.58** | 30.99** | 28.96** | 89.19** | 18.40** | — | 4.99* | 3.06 | 13.20** | 0.96 | 0.85 | 5.33* | 5.84* | 20.22** | 5.08* | 42.11** | 11.61** | 29.94** | 6.74** | 7.38** | 0.47 | |
| SB | 8.77** | 13.77** | 10.27** | 67.10** | 8.13** | 4.99* | — | 10.69** | 4.56* | 0.25 | 0.30 | 0.73 | 15.08** | 38.23** | 0.48 | 23.46** | 2.41 | 14.00** | 1.43 | 20.55** | 1.47 | |
| Llama | T | 34.18** | 35.79** | 30.30** | 89.28** | 28.34** | 3.06 | 10.69** | — | 37.09** | 9.03** | 8.62** | 16.12** | 0.35 | 6.28* | 19.19** | 58.08** | 23.16** | 43.91** | 18.64** | 0.02 | 5.61* |
| TS | 0.21 | 0.69 | 0.06 | 22.45** | 0.03 | 13.20** | 4.56* | 37.09** | — | 7.81** | 8.57** | 1.61 | 35.74** | 59.09** | 1.75 | 5.46* | 0.56 | 1.20 | 0.64 | 32.53** | 9.71** | |
| TB | 11.05** | 12.10** | 10.25** | 59.74** | 7.84** | 0.96 | 0.25 | 9.03** | 7.81** | — | 0.00 | 1.69 | 15.84** | 36.84** | 1.71 | 26.06** | 4.14* | 16.75** | 2.70 | 10.61** | 0.16 | |
| TSB | 10.75** | 12.30** | 10.16** | 58.72** | 8.07** | 0.85 | 0.30 | 8.62** | 8.57** | 0.00 | — | 1.92 | 15.09** | 35.27** | 1.83 | 26.38** | 4.38* | 17.32** | 2.98 | 9.67** | 0.11 | |
| S | 2.56 | 4.69* | 2.13 | 33.34** | 3.54 | 5.33* | 0.73 | 16.12** | 1.61 | 1.69 | 1.92 | — | 23.06** | 45.22** | 0.00 | 14.49** | 0.18 | 5.32* | 0.20 | 17.86** | 3.01 | |
| B | 39.32** | 41.32** | 37.80** | 96.75** | 34.04** | 5.84* | 15.08** | 0.35 | 35.74** | 15.84** | 15.09** | 23.06** | — | 5.76* | 21.04** | 69.73** | 31.29** | 53.60** | 25.26** | 0.14 | 10.29** | |
| SB | 62.01** | 69.72** | 65.91** | 132.08** | 65.24** | 20.22** | 38.23** | 6.28* | 59.09** | 36.84** | 35.27** | 45.22** | 5.76* | — | 38.22** | 92.66** | 51.61** | 76.55** | 48.80** | 6.19* | 30.08** | |
| Mistral | T | 4.50* | 4.72* | 2.96 | 41.44** | 2.02 | 5.08* | 0.48 | 19.19** | 1.75 | 1.71 | 1.83 | 0.00 | 21.04** | 38.22** | — | 18.21** | 0.43 | 7.90** | 0.17 | 19.38** | 2.88 |
| TS | 4.21* | 2.79 | 5.33* | 7.87** | 5.47* | 42.11** | 23.46** | 58.08** | 5.46* | 26.06** | 26.38** | 14.49** | 69.73** | 92.66** | 18.21** | — | 13.56** | 2.79 | 11.83** | 69.98** | 32.90** | |
| TB | 2.34 | 3.31 | 1.70 | 45.75** | 0.98 | 11.61** | 2.41 | 23.16** | 0.56 | 4.14* | 4.38* | 0.18 | 31.29** | 51.61** | 0.43 | 13.56** | — | 9.75** | 0.00 | 43.68** | 8.44** | |
| TSB | 0.39 | 0.04 | 0.88 | 22.90** | 0.76 | 29.94** | 14.00** | 43.91** | 1.20 | 16.75** | 17.32** | 5.32* | 53.60** | 76.55** | 7.90** | 2.79 | 9.75** | — | 3.62 | 68.73** | 24.89** | |
| S | 1.46 | 2.77 | 1.07 | 29.36** | 1.90 | 6.74** | 1.43 | 18.64** | 0.64 | 2.70 | 2.98 | 0.20 | 25.26** | 48.80** | 0.17 | 11.83** | 0.00 | 3.62 | — | 20.30** | 4.24* | |
| B | 46.89** | 53.35** | 54.30** | 119.71** | 38.75** | 7.38** | 20.55** | 0.02 | 32.53** | 10.61** | 9.67** | 17.86** | 0.14 | 6.19* | 19.38** | 69.98** | 43.68** | 68.73** | 20.30** | — | 20.33** | |
| SB | 15.42** | 20.66** | 18.06** | 72.90** | 12.04** | 0.47 | 1.47 | 5.61* | 9.71** | 0.16 | 0.11 | 3.01 | 10.29** | 30.08** | 2.88 | 32.90** | 8.44** | 24.89** | 4.24* | 20.33** | — | |
| GPT | Llama | Mistral | ||||||||||||||||||||
| T | TS | TB | TSB | S | B | SB | T | TS | TB | TSB | S | B | SB | T | TS | TB | TSB | S | B | SB | ||
| GPT | T | — | 34.40** | 20.11** | 449.51** | 341.98** | 188.99** | 377.91** | 96.77** | 32.68** | 38.99** | 224.78** | 66.33** | 47.50** | 172.07** | 51.50** | 2.29 | 18.13** | 288.68** | 232.80** | 135.35** | 287.96** |
| TS | 34.40** | — | 6.67** | 435.86** | 328.99** | 158.38** | 362.25** | 200.31** | 132.25** | 140.91** | 218.99** | 60.72** | 42.21** | 163.93** | 136.48** | 19.91** | 82.16** | 265.40** | 231.29** | 113.12** | 264.55** | |
| TB | 20.11** | 6.67** | — | 440.16** | 344.31** | 173.22** | 374.56** | 163.22** | 81.73** | 103.88** | 233.55** | 68.50** | 50.77** | 176.20** | 104.43** | 3.12 | 62.26** | 283.78** | 240.87** | 128.75** | 283.02** | |
| TSB | 259.15** | 373.77** | 318.02** | — | 24.89** | 175.03** | 15.45** | 68.03** | 125.03** | 124.95** | 107.43** | 249.53** | 286.51** | 146.23** | 130.00** | 286.31** | 175.39** | 83.12** | 75.74** | 204.39** | 74.29** | |
| S | 253.85** | 372.88** | 325.06** | 24.89** | — | 61.13** | 4.88* | 73.52** | 136.62** | 132.00** | 36.33** | 186.29** | 197.09** | 72.16** | 134.29** | 293.99** | 177.19** | 12.26** | 33.39** | 93.04** | 14.84** | |
| B | 199.10** | 272.44** | 239.34** | 175.03** | 61.13** | — | 91.51** | 28.50** | 61.03** | 65.13** | 7.92** | 20.40** | 40.57** | 0.19 | 75.61** | 192.91** | 120.69** | 33.56** | 13.44** | 8.85** | 33.00** | |
| SB | 245.87** | 363.62** | 310.64** | 15.45** | 4.88* | 91.51** | — | 64.08** | 121.96** | 118.26** | 51.81** | 201.48** | 222.84** | 96.82** | 127.00** | 283.93** | 168.95** | 24.57** | 44.52** | 124.66** | 31.57** | |
| Llama | T | 96.77** | 200.31** | 163.22** | 433.77** | 350.42** | 149.02** | 375.46** | — | 21.99** | 21.16** | 236.90** | 72.80** | 54.29** | 180.28** | 18.25** | 138.91** | 40.15** | 276.51** | 263.02** | 114.96** | 266.27** |
| TS | 32.68** | 132.25** | 81.73** | 475.86** | 395.04** | 164.84** | 413.03** | 21.99** | — | 0.00 | 274.00** | 103.51** | 75.98** | 215.70** | 0.00 | 85.23** | 5.80* | 299.13** | 317.34** | 131.29** | 295.40** | |
| TB | 38.99** | 140.91** | 103.88** | 468.49** | 379.28** | 163.60** | 400.96** | 21.16** | 0.00 | — | 271.24** | 99.85** | 76.18** | 220.90** | 0.00 | 81.82** | 8.23** | 299.39** | 307.37** | 133.30** | 297.66** | |
| TSB | 252.24** | 383.38** | 327.86** | 107.43** | 36.33** | 7.92** | 51.81** | 79.93** | 148.70** | 144.11** | — | 78.32** | 130.33** | 12.27** | 136.57** | 313.88** | 184.20** | 7.11** | 2.32 | 28.24** | 5.32* | |
| S | 262.93** | 399.75** | 341.51** | 249.53** | 186.29** | 20.40** | 201.48** | 96.50** | 180.50** | 174.87** | 78.32** | — | 2.49 | 46.08** | 160.35** | 341.84** | 201.12** | 102.45** | 121.28** | 7.80** | 103.54** | |
| B | 260.62** | 393.63** | 345.41** | 286.51** | 197.09** | 40.57** | 222.84** | 107.51** | 190.68** | 185.48** | 130.33** | 2.49 | — | 88.67** | 164.74** | 343.47** | 215.15** | 147.84** | 126.48** | 23.21** | 135.88** | |
| SB | 265.39** | 390.41** | 337.63** | 146.23** | 72.16** | 0.19 | 96.82** | 89.64** | 161.96** | 160.99** | 12.27** | 46.08** | 88.67** | — | 153.62** | 325.90** | 201.51** | 27.24** | 18.31** | 8.15** | 27.36** | |
| Mistral | T | 51.50** | 136.48** | 104.43** | 451.99** | 360.36** | 167.31** | 400.83** | 18.25** | 0.00 | 0.00 | 237.93** | 79.88** | 57.59** | 186.23** | — | 86.05** | 9.88** | 282.52** | 272.88** | 131.00** | 295.80** |
| TS | 2.29 | 19.91** | 3.12 | 465.55** | 366.27** | 173.09** | 401.68** | 138.91** | 85.23** | 81.82** | 249.35** | 84.57** | 60.26** | 194.64** | 86.05** | — | 33.00** | 295.92** | 276.45** | 131.27** | 291.16** | |
| TB | 18.13** | 82.16** | 62.26** | 453.61** | 360.22** | 187.09** | 393.63** | 40.15** | 5.80* | 8.23** | 240.67** | 76.52** | 59.10** | 191.91** | 9.88** | 33.00** | — | 295.82** | 255.13** | 143.73** | 299.24** | |
| TSB | 240.52** | 343.16** | 295.33** | 83.12** | 12.26** | 33.56** | 24.57** | 56.78** | 104.94** | 107.75** | 7.11** | 102.45** | 147.84** | 27.24** | 113.84** | 263.56** | 161.16** | — | 0.71 | 74.25** | 0.02 | |
| S | 270.72** | 411.28** | 350.63** | 75.74** | 33.39** | 13.44** | 44.52** | 97.52** | 180.09** | 169.62** | 2.32 | 121.28** | 126.48** | 18.31** | 163.77** | 345.38** | 205.33** | 0.71 | — | 32.63** | 0.51 | |
| B | 195.26** | 276.06** | 247.61** | 204.39** | 93.04** | 8.85** | 124.66** | 29.48** | 64.62** | 69.14** | 28.24** | 7.80** | 23.21** | 8.15** | 77.65** | 203.68** | 125.25** | 74.25** | 32.63** | — | 77.30** | |
| SB | 219.47** | 317.90** | 275.13** | 74.29** | 14.84** | 33.00** | 31.57** | 47.58** | 94.93** | 95.66** | 5.32* | 103.54** | 135.88** | 27.36** | 104.22** | 243.56** | 146.45** | 0.02 | 0.51 | 77.30** | — | |
| GPT | Llama | Mistral | ||||||||||||||||||||
| T | TS | TB | TSB | S | B | SB | T | TS | TB | TSB | S | B | SB | T | TS | TB | TSB | S | B | SB | ||
| GPT | T | — | 277.85** | 32.27** | 255.80** | 437.95** | 16.75** | 383.38** | 23.89** | 60.47** | 6.33* | 25.05** | 280.54** | 48.25** | 141.66** | 17.31** | 186.27** | 47.23** | 33.34** | 339.27** | 106.59** | 99.26** |
| TS | 277.85** | — | 195.62** | 3.90* | 139.76** | 316.95** | 94.82** | 258.18** | 26.08** | 212.25** | 81.25** | 28.02** | 301.64** | 3.96* | 272.09** | 4.23* | 357.08** | 139.18** | 62.00** | 448.55** | 43.41** | |
| TB | 32.27** | 195.62** | — | 192.83** | 359.77** | 74.74** | 311.69** | 56.81** | 26.50** | 30.13** | 3.88* | 212.64** | 91.74** | 84.88** | 52.94** | 118.53** | 107.43** | 4.20* | 262.49** | 184.46** | 49.97** | |
| TSB | 255.80** | 3.90* | 192.83** | — | 150.75** | 318.84** | 109.80** | 240.60** | 17.80** | 200.62** | 70.24** | 38.68** | 290.01** | 0.96 | 253.02** | 0.56 | 342.79** | 124.91** | 73.86** | 437.54** | 31.74** | |
| S | 437.95** | 139.76** | 359.77** | 150.75** | — | 518.92** | 27.23** | 483.85** | 187.55** | 434.56** | 284.06** | 20.53** | 540.53** | 150.78** | 481.42** | 149.24** | 547.39** | 324.18** | 57.02** | 656.50** | 263.31** | |
| B | 16.75** | 316.95** | 74.74** | 318.84** | 518.92** | — | 480.17** | 2.13 | 97.45** | 0.63 | 57.42** | 381.03** | 15.78** | 215.47** | 0.13 | 239.73** | 6.63* | 81.30** | 425.04** | 67.27** | 202.21** | |
| SB | 383.38** | 94.82** | 311.69** | 109.80** | 27.23** | 480.17** | — | 441.35** | 150.71** | 393.49** | 243.99** | 7.00** | 508.99** | 119.71** | 437.53** | 109.73** | 505.97** | 279.39** | 21.33** | 619.11** | 229.59** | |
| Llama | T | 23.89** | 258.18** | 56.81** | 240.60** | 483.85** | 2.13 | 441.35** | — | 247.06** | 12.31** | 132.97** | 427.63** | 7.13** | 249.42** | 2.17 | 260.91** | 0.13 | 80.77** | 412.95** | 14.41** | 145.35** |
| TS | 60.47** | 26.08** | 26.50** | 17.80** | 187.55** | 97.45** | 150.71** | 247.06** | — | 164.01** | 24.63** | 106.58** | 224.01** | 12.95** | 162.56** | 17.64** | 151.72** | 16.56** | 119.78** | 203.03** | 0.07 | |
[…]

### v3

@@ Appendix C Sources can be misleading
Table 8 reports the accuracy of the LLMs on these specific test cases for liberal and conservative sources respectively. We see that the LLMs suffer from incorrect classifications when only the source is considered as the ground truth contradicts the ideology of the source. Here, providing the title increases the accuracy of the model especially in the conservative sources where the accuracy on just the source is for the YouTube and Ad Fontes datasets. In fact, providing the title for the YouTube test set in this case still does not improve the accuracy seeing as the model accuracy is still . This suggests that the LLM overtly relies on the source in this case completely disregarding title wherein accuracy on just the title was close to . Similar trends can be seen for the Ad Fontes news test set where the accuracy of source alone was and title was , and both was . We observe that the AllSides news test set still has better performance when the source is provided at a . On closer inspection, we observe that in the test set sampled, there is a high correlation between the source and article ideology which likely leads to such a high performance when source is provided.
Appendix D Few Shot Results for all Datasets
Figure 4 shows the increases in accuracy across all datasets by increasing the number of shots.
Appendix E Ablations on different parameters
E.1 MLP Baseline
| YouTube | Ad Fontes | AllSides | |||
| Title | Source | Body | Slant | News | News |
| ✓ | 0.51 | 0.38 | 0.43 | ||
| ✓ | 0.61 | 0.55 | 0.69 | ||
| ✓ | 0.55 | 0.38 | 0.47 | ||
| ✓ | ✓ | 0.63 | 0.62 | 0.57 | |
| ✓ | ✓ | 0.57 | 0.43 | 0.46 | |
| ✓ | ✓ | 0.62 | 0.60 | 0.58 | |
| ✓ | ✓ | ✓ | 0.64 | 0.61 | 0.52 |
E.2 RoBERTa baseline
| YouTube | Ad Fontes | AllSides | |||
| Title | Source | Body | Slant | News | News |
| ✓ | 0.57 | 0.35 | 0.52 | ||
| ✓ | 0.45 | 0.48 | 0.48 | ||
| ✓ | 0.62 | 0.35 | 0.50 | ||
| ✓ | ✓ | 0.65 | 0.44 | 0.58 | |
| ✓ | ✓ | 0.64 | 0.38 | 0.53 | |
| ✓ | ✓ | 0.65 | 0.44 | 0.58 | |
| ✓ | ✓ | ✓ | 0.65 | 0.44 | 0.55 |
E.3 GPT-4o
Table 11 shows the accuracy of the GPT-4o model under different ablations of title, source, and description across the three datasets and varying number of demonstrations considered. We see that the model performs the best when given a higher number of demonstrations for the YouTube slant and Ad Fontes news datasets. For the YouTube dataset, the model performed best when title, source, and description were all provided whereas for the Ad Fontes news dataset, only title and source were sufficient. The AllSides news dataset showed a different trend in that zero demonstrations gave the best score when only the source was provided. We observed in the test set for this dataset that source and target label were highly correlated (i.e., a video from Fox, a well-known conservative outlet, was almost always labeled conservative) making it simpler for the model to just assign the source ideology as the target ideology. In Appendix C, we investigate this further by considering edge cases in the test set.
| YouTube | Ad Fontes | AllSides | ||||
| Title | Source | Body | Slant | News | News | |
| 0 | ✓ | 0.64 | 0.69 | 0.50 | ||
| 0 | ✓ | 0.63 | 0.80 | 0.88 | ||
| 0 | ✓ | 0.56 | 0.64 | 0.44 | ||
| 0 | ✓ | ✓ | 0.64 | 0.78 | 0.74 | |
| 0 | ✓ | ✓ | 0.63 | 0.75 | 0.55 | |
| 0 | ✓ | ✓ | 0.59 | 0.82 | 0.85 | |
| 0 | ✓ | ✓ | ✓ | 0.71 | 0.86 | 0.72 |
| 4 | ✓ | 0.67 | 0.74 | 0.55 | ||
| 4 | ✓ | ✓ | 0.73 | 0.87 | 0.83 | |
| 4 | ✓ | ✓ | 0.69 | 0.80 | 0.58 | |
| 4 | ✓ | ✓ | ✓ | 0.72 | 0.86 | 0.82 |
| 8 | ✓ | 0.68 | 0.73 | 0.55 | ||
| 8 | ✓ | ✓ | 0.72 | 0.87 | 0.86 | |
| 8 | ✓ | ✓ | 0.70 | 0.81 | 0.59 | |
| 8 | ✓ | ✓ | ✓ | 0.71 | 0.86 | 0.84 |
| 12 | ✓ | 0.69 | 0.74 | 0.55 | ||
| 12 | ✓ | ✓ | 0.74 | 0.87 | 0.87 | |
| 12 | ✓ | ✓ | 0.71 | 0.82 | 0.59 | |
| 12 | ✓ | ✓ | ✓ | 0.71 | 0.87 | 0.84 |
E.4 Llama2
| YouTube | Ad Fontes | AllSides | ||||
| Title | Source | Body | Slant | News | News | |
| 0 | ✓ | 0.53 | 0.49 | 0.42 | ||
| 0 | ✓ | 0.61 | 0.52 | 0.82 | ||
| 0 | ✓ | 0.52 | 0.49 | 0.37 | ||
| 0 | ✓ | ✓ | 0.63 | 0.57 | 0.65 | |
| 0 | ✓ | ✓ | 0.58 | 0.57 | 0.46 | |
| 0 | ✓ | ✓ | 0.48 | 0.65 | 0.71 | |
| 0 | ✓ | ✓ | ✓ | 0.58 | 0.69 | 0.59 |
| 4 | ✓ | 0.64 | 0.60 | 0.54 | ||
| 4 | ✓ | ✓ | 0.68 | 0.79 | 0.75 | |
| 4 | ✓ | ✓ | 0.66 | 0.68 | 0.51 | |
| 4 | ✓ | ✓ | ✓ | 0.64 | 0.79 | 0.68 |
| 8 | ✓ | 0.60 | 0.56 | 0.51 | ||
| 8 | ✓ | ✓ | 0.63 | 0.76 | 0.74 | |
| 8 | ✓ | ✓ | 0.63 | 0.67 | 0.52 | |
| 8 | ✓ | ✓ | ✓ | 0.60 | 0.78 | 0.67 |
| 12 | ✓ | 0.60 | 0.60 | 0.50 | ||
| 12 | ✓ | ✓ | 0.62 | 0.75 | 0.72 | |
| 12 | ✓ | ✓ | 0.58 | 0.68 | 0.49 | |
| 12 | ✓ | ✓ | ✓ | 0.56 | 0.78 | 0.67 |
E.5 Mistral
| YouTube | Ad Fontes | AllSides | ||||
| Title | Source | Body | Slant | News | News | |
| 0 | ✓ | 0.61 | 0.57 | 0.44 | ||
| 0 | ✓ | 0.61 | 0.72 | 0.83 | ||
| 0 | ✓ | 0.52 | 0.59 | 0.34 | ||
| 0 | ✓ | ✓ | 0.67 | 0.72 | 0.71 | |
| 0 | ✓ | ✓ | 0.62 | 0.62 | 0.41 | |
| 0 | ✓ | ✓ | 0.57 | 0.73 | 0.66 | |
| 0 | ✓ | ✓ | ✓ | 0.65 | 0.74 | 0.58 |
| 4 | ✓ | 0.61 | 0.57 | 0.42 | ||
| 4 | ✓ | ✓ | 0.65 | 0.73 | 0.73 | |
| 4 | ✓ | ✓ | 0.62 | 0.61 | 0.44 | |
| 4 | ✓ | ✓ | ✓ | 0.63 | 0.68 | 0.60 |
| 8 | ✓ | 0.62 | 0.59 | 0.44 | ||
| 8 | ✓ | ✓ | 0.64 | 0.71 | 0.73 | |
| 8 | ✓ | ✓ | 0.63 | 0.62 | 0.48 | |
| 8 | ✓ | ✓ | ✓ | 0.63 | 0.68 | 0.61 |
| 12 | ✓ | 0.64 | 0.61 | 0.44 | ||
| 12 | ✓ | ✓ | 0.66 | 0.71 | 0.73 | |
| 12 | ✓ | ✓ | 0.61 | 0.63 | 0.48 | |
| 12 | ✓ | ✓ | ✓ | 0.63 | 0.69 | 0.61 |
Appendix F Candidate set size can be smaller
| LLM | YouTube Slant | Ad Fontes News | AllSides News |
| GPT-4o | 0.66 | 0.74 | 0.52 |
| Llama2 | 0.60 | 0.57 | 0.50 |
| Mistral | 0.63 | 0.59 | 0.42 |
Appendix G Balanced demonstration selection is better than random
| Dataset | Random | Balanced BERTScore |
| YouTube Slant | 0.52 | 0.61 |
| Ad Fontes News | 0.49 | 0.57 |
| AllSides News | 0.39 | 0.42 |
Appendix H Qualitative results for the political news analysis for YouTube videos
Qualitative results for political news videos, non-political news videos, and political non-news videos that changed predictions as a result of adding information about the source are provided in Tables 16, 17, and 18 respectively.
| Title | Channel |
| Liberal to Neutral | |
| How welcome are Russian deserters in Georgia? | Focus on Europe | DW News |
| Ukrainians ditch Russian traditions and celebrate Christmas with the West | Sky News |
| Asian Americans Debate Model Minority & Asian Hate | VICE Debates | VICE |
| Secret Service under scrutiny ahead January 6 committee’s final report | PBS NewsHour |
| White House Correspondent Jeff Mason on his tense exchanges with Trump | Reuters |
| Liberal to Conservative | |
| Democrats introduce bill to block Trump from holding office | Newsmax |
| LIVE: Jan. 6 Committee Votes on Criminal Referrals, Final Report | NTD |
| New CRIMINAL REFERRALS Against Former President Trump Unveiled by Jan. 6th Committee | BlazeTV |
| Tonight’s Tipping Points: Hate Crimes, Planned Parenthood, & Trans Military Members! | One America News Network |
| Trolling #AMFEST2022 Attendees as Left-Winger ’Garth Griffin’ of Edge News | Project Veritas |
| Neutral to Liberal | |
| Haberman says Trump has been diminished but isn’t dead politically | CNN |
| Kevin McCarthy’s Path To Speaker Is In Peril | MSNBC |
| Anger grows in Russia over Ukraine’s New Year’s Day strike on Russian troops | CNN |
| Rep. Adam Kinzinger Delivers Final Speech As Congressman | MSNBC |
| ’Shocked’: Russia expert explains why Putin’s change of word is significant | CNN |
| Neutral to Conservative | |
| UK Weather: What’s behind Britain’s big freeze? | UK snow chaos explained | Daily Mail |
| Transgender Clinical Psychologist on people de-transitioning | Dr. Erica Anderson | Newsmax |
| Actor slams crypto as ’biggest Ponzi scheme ever’ during Senate hearing | Fox News |
| Rep. Spartz explains her reasoning for voting ’present’ for Speaker | Fox News |
| Internet users allege Bryan Kohberger attended Idaho victims’ vigil | Fox News |
| Conservative to Liberal | |
| Trump Supporter: Too Much Diversity Will Kill America | The Young Turks |
| Russian foreign minister: Ukraine must give up occupied territories | CNN |
| These Trump Supporters Are CONVINCED "Sleepy Joe Poopy Pants Isn’t The Acting President" | The Young Turks |
| Steve Bannon reacts to Trump’s ’major announcement’ | CNN |
| Trump adviser tells reporter ’the magic is gone’ in new article about re-election bid | CNN |
| Conservative to Neutral | |
| Justin Amash APPLAUDS Revolt Against McCarthy, Says Ruling Oligarchy CRUSHES Dissent | The Hill |
| State Of Emergency Declared At Texas Border | NBC News |
| Watch Trump claim the army retook airports during the Revolutionary War in July Fourth speech | The Globe and Mail |
| Trump Calls Jan. 6 Committee A ’Kangaroo Court’ After Criminal Referrals | NBC News |
| ’They shot up the WRONG house!’ Sheriff Grady Judd on young kids shot at while in bed in Polk County | WFLA News Channel 8 |
| Title | Channel |
| Liberal to Neutral | |
| TRIPLEDEMIC Or Free Market Failure To Blame For Medicine SHORTAGE? STRUGGLE To Find Children’s Meds | The Hill |
| Watch: Jan. 6 Committee Presents Key Video Evidence Compilation | NBC News |
| Neutral to Liberal | |
| Top U.S. & World Headlines — December 27, 2022 | Democracy Now! |
| Top U.S. & World Headlines — January 3, 2023 | Democracy Now! |
| Raskin On Cancer Diagnosis: ’Totally My Plan To Make It Through This Thing’ | MSNBC |
| Watch The 11th Hour Highlights: Dec. 20 | MSNBC |
| Plans For Zelenskyy D.C. Trip Came Together In A ’Matter Of Days’ | MSNBC |
| Neutral to Conservative | |
| McCarthy speakership battle: What happens if he concedes? | Fox Business |
| The yule log Christmas tradition | Fox News |
| “Woman” Is Named Word of the Year | DailyWire+ |
| What is the controversy surrounding Meghan Markle and her sister? | Fox News |
| ’DOUBLE WHAMMY’: Real estate expert issues pricing warning | Fox Business |
| Conservative to Liberal | |
| All Gestures, All Shock Opera: DeSantis Launches His Newest Battle | MSNBC |
| Christian Influencer Makes Surprising Confession About Trans Porn | The Young Turks |
| The War on Christmas RAGES ON! | The Young Turks |
| Alex Jones Says He’s Paying His Personal Trainer $100,000 A Week | The Young Turks |
| ’Look at where he is today’: Bash on Kinzinger’s farewell speech | CNN |
| Conservative to Neutral | |
| Cancelling the F-22 Fighter is Worse than You Think | Task & Purpose |
| Title | Channel |
| Liberal to Neutral | |
| Is capitalism actually broken? | TED-Ed |
| Update from Ukraine | Ukraine is getting ready for the winter Attack | Ruzzia will lose it | Denys Davydov |
| Van Lathan: A Jewish Producer Confronted Kanye at TMZ when Ye Said He Loves Hitler (Part 2) | djvlad |
| The Trouble with the Electoral College | CGP Grey |
| Those Times Sam Harris Made Us Slap Our Knees | Enlightainment |
| Liberal to Conservative | |
| The New Beginning of Same-Sex Marriage | Ben Shapiro |
| Meet Kamala Harris’s 6-Year-Old Speechwriter | The Babylon Bee |
| Pro-Choice Bosses - A Generous Offer To Our Pregnant Employees | Live Action |
| Confronting Charlie Kirk and Tim Pool at Turning Point USA America Fest | Alex Stein |
| I Showed Up To An Anti-Gun Protest In Los Angeles | Let’s Talk | James Klug |
| Neutral to Liberal | |
| Meanwhile… Pantone’s Color of the Year | Macron Makes Condoms Free In France | The Late Show with Stephen Colbert |
| Grandma Arrested For Feeding The Homeless?! Danbury Court Case Clarifcation + GIVEAWAY! | Long Island Audit |
| LIVE: EMERGENCY January 6 Committee hearing - Day 6 | Brian Tyler Cohen |
| The Funniest Donald Trump Stories On The Graham Norton Show | Part One | The Graham Norton Show |
| Prison Officer BUSTED After Massive Nazi Gang Attack Caught | The Damage Report |
| Neutral to Conservative | |
| Winning With The Hand You’re Dealt with Rick Warren | Saddleback Church |
| Fighting the Darkest Criminal Enterprise on the Planet | The 700 Club |
| STORE WARS: Target, Walmart & Home Depot go to war against Organized Retail theft | News For Reasonable People |
| Warner MAY BE SUED over Amber Heard and Aquaman 2?! NEW Hollywood Lawsuit! | ThatUmbrellaGuy |
| Historical alteration on the Ukraine Map: Russian 1st Corps withdrawing en masse from 4 locations! | Divine Justice |
| Conservative to Liberal | |
| Elon Musk Finally CRACKS In Ultimate Outburst Over Free Speech | The Damage Report |
| Chris Wallace: This Would Never Have Happened Under Nancy Pelosi | The Late Show with Stephen Colbert |
| Mike Lindell Snaps After Hannity Admits There Was No Fraud | Tommy Campbell |
| Trump ATTACKS the FBI and DOJ and Loses his Mind in NEW POSTS | MeidasTouch |
| BREAKING: Kari Lake slammed with brutal legal news | Brian Tyler Cohen |
| Conservative to Neutral | |
| Why US Let Russia Become the Biggest Nuclear Power | The Infographics Show |
| The CIA’s Biggest Scam Exposed | KiraTV |
| Cop Pulls Over "Legal Expert" And SCHOOLS HIM On The Law | Audit the Audit |
| Why US Enemies are Scared of Ghosts (Elite Special Forces) | The Infographics Show |
| "INSPIRATIONAL HOLIDAY VIDEO" — A Bad Lip Reading of Joe Biden | Bad Lip Reading |
Appendix I Placement of popular news channels based on the slant cut-offs
Figure 5 shows the placement of popular news channels based on the YouTube slant dataset. Corresponding cut-offs at -0.33 and +0.33 are also shown for the liberal and conservative respectively.
Appendix J Complete set of results for the various LLM configurations
Table 19 shows the final set of results for all configurations.
| Metadata | Dataset | ||||||
| Model | Shots | Title | Source | Body | YouTube Dataset | Ad Fontes News | AllSides News |
| MLP | — | ✓ | 0.49 | 0.38 | 0.41 | ||
| RoBERTa | — | ✓ | 0.65 | 0.33 | 0.57 | ||
| Baly et al. (2020) | — | ✓ | — | — | 0.46 | ||
| GPT-4o | 0 | ✓ | 0.64 | 0.69 | 0.50 | ||
| Llama2-13B | 0 | ✓ | 0.44 | 0.46 | 0.40 | ||
| Mistral-7B | 0 | ✓ | 0.59 | 0.55 | 0.44 | ||
| GPT-4o | 0 | ✓ | ✓ | 0.52 | 0.78 | 0.74 | |
| Llama2-13B | 0 | ✓ | ✓ | 0.43 | 0.54 | 0.64 | |
| Mistral-7B | 0 | ✓ | ✓ | 0.67 | 0.72 | 0.71 | |
| GPT-4o | 0 | ✓ | ✓ | 0.57 | 0.75 | 0.55 | |
| Llama2-13B | 0 | ✓ | ✓ | 0.49 | 0.55 | 0.46 | |
| Mistral-7B | 0 | ✓ | ✓ | 0.59 | 0.62 | 0.41 | |
[…]